# Bimanual Deformable Bag Manipulation Using a Structure-of-Interest Based Latent Dynamics Model

### v1

@@ IV-C Model Predictive Control
Figure 6: Experimental set-up. (a) Overview of the experimental set-up to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks. (b)-(d) The SOI shape categories are considered in experiments: Long Oval (LO), Round Oval (RO), and Short Oval (SO).
TABLE I: Averaged SOI particle sampling results across 90 frames.
@@
V Experiments
V-A Experiment Setup
As shown in Fig. 6, we present the general experimental setup used to validate our SOI-based dynamics model for bimanual deformable bag manipulation tasks. The setup includes four RealSense D435 RGBD cameras, each positioned at a corner, to capture RGBD images of the deformable fabric bag from multiple angles at a frequency of 30Hz and a resolution of 640$\times$480. Two Dobot CR5 robotic manipulators, equipped with 7 degrees of freedom, grasp the bag’s handles using 3D-printed grippers secured with zip ties. The Structure of Interest (SOI) of the bag, specifically the opening rim, has been modified for better SOI perception: it is cut and sewn with the same type of fabric but in a contrasting green color, which facilitates the detection while maintaining the consistency of the bag’s dynamic behavior. Additionally, a top-down camera and a front-facing camera are employed to capture the experimental process from various perspectives.
A dataset of 18,000 frames, around 30 minutes, comprises the training data, featuring 200 episodes, each spanning 90 frames. Within every episode, we execute six actions on the fabric bag. Before each episode’s inception, the fabric bag is orderly arranged to ensure the opening rim protrudes outward without excessive wrinkling. The data collection policy indiscriminately selects from the action space parameters after doing collision checking and bag constraint checking, which include translations along the $x$, $y$, $z$ axes, and the rotation around $rz$. Throughout each episode, we archive partial point clouds captured by RGB-D cameras, as well as the robot end-effector’s poses as returned by the robot controller. Data recording is contingent on the gripper’s movement, an approach taken to optimize memory consumption.
As illustrated in Fig. 6(b)-(d), this study considers three SOI shape categories: Long Oval (LO), Round Oval (RO), and Short Oval (SO). The experiment comprises two types of manipulation tasks. In the SOI shape preservation task, the robotic manipulators aim to keep the SOI shape consistent while moving the fabric bag from its initial position to a predefined target location. Both the initial and target configurations share the same shape but differ in their spatial transformations. Conversely, the SOI shape servoing task involves altering the SOI shape from one configuration to another distinct shape configuration by manipulating the bag handles. We conduct a quantitative evaluation of each component within our framework, encompassing SOI particle sampling, the SOI dynamics model, and performance results of designed manipulation tasks. Our primary metrics for evaluating the manipulation performance are Chamfer distance (CD) and Earth mover’s distance (EMD), along with a hybrid metric that combines both. Furthermore, we employ Geodesic distance (GD) [20] to assess the proximity of boundary points residing on one manifold with boundary to those on another.
@@ V-A Experiment Setup
TABLE II: The performance of the SOI particle dynamics model with different loss function.
@@
Figure 7: Qualitative results in (a) SOI shape preserving and (b) SOI shape servoing experiments.
V-B SOI Particle Sampling
We commence by benchmarking the proposed Global Particle Sampling (GPS) technique against the Local Particle Sampling (LPS) baseline. LPS initiates by processing the SOI-specific partial point cloud of the fabric bag through its unique color signature. Subsequently, it encapsulates the incomplete SOI-specific cloud with a convex hull. Following this, it performs the point sampling procedure, eventually amalgamating the sampled points into a full set of SOI-specific particles. As shown in Table I, we compute the mean distance between the sampled particles and the ground-truth particles, the latter captured by a professional 3D scanner. Our analysis indicates that the GPS method incurs lower losses in terms of all distance metrics, thus outperforming the LPS approach. These outcomes align with the premise that incorporating additional topological information about the SOI can markedly enhance the quality of sampling, particularly in scenarios where occlusions are present.
V-C GNN-based SOI Dynamics Model
The training of our GNN model, detailed in Section IV-B, commences with the construction of a graph, where edges are formed between vertices that are within a distance threshold of $d=0.04$. Every vertex and edge are encoded using 3-layer Multilayer Perceptrons (MLPs), featuring hidden and output layers, each with 300 neurons. The propagation module is constituted by a fully connected neural layer with a layer size of 300. For motion prediction, we employ an additional 3-layer MLP with the hidden layer configured to 300 neurons. ReLU activation functions are utilized throughout the neural networks to introduce non-linearity. The model undergoes training for 120 epochs, employing the Adam optimizer. We have selected a batch size of 32 to balance the trade-off between generalization and computational efficiency. The learning rate is set at $5\mathrm{e}$-4, which is a conventional choice for steady convergence. These hyperparameters were chosen to foster a robust learning process while maintaining the capacity to capture complex patterns within the data.
Subsequently, we evaluate the performance of GNN-based SOI dynamics models using different loss functions. Table 6 indicates that only using CD or EMD individually optimizes for its own metric but not others, but combining CD and EMD losses leads to better overall performance on all metrics, and the Geodesic distance is significantly improved, compared to using them individually. The combination of $0.85~{}\text{CD}+0.15~{}\text{EMD}$ achieves the best performance across all metrics, giving the lowest CD, competitive EMD, and lowest GD, so we select this hybrid distance for the following experiments. The reason is that CD and EMD losses capture different aspects of point cloud similarity. CD compares point-wise distances while EMD measures global shape differences. By combining them, the model is optimized for both local point accuracy and global shape matching. The proper combination of 0.85 CD + 0.15 EMD balances these objectives best.
TABLE III: Mean Hybrid Distance Error for SOI Shape Preserving by Different Dynamics Modeling Methods
@@ V-C GNN-based SOI Dynamics Model
TABLE IV: Mean Hybrid Distance Error and Success Rate for SOI Shape Servoing by Different Methods
Method Mean Hybrid Distance Error Success Rate
LO $\rightarrow$ SO SO $\rightarrow$ RO LO $\rightarrow$ RO LO $\rightarrow$ SO SO $\rightarrow$ RO LO $\rightarrow$ RO Total
VS [33] $3.75\pm 0.19$ $3.69\pm 0.22$ $3.92\pm 0.23$ $21/30$ $19/30$ $22/30$ $68.89\%$
LSC [34] $2.94\pm 0.48$ $3.14\pm 0.42$ $2.89\pm 0.44$ $24/30$ $24/30$ $26/30$ $82.23\%$
MSM [35] $3.22\pm 0.43$ $3.49\pm 0.47$ $3.28\pm 0.39$ $19/30$ $21/30$ $23/30$ $70.00\%$
FEM [36] $2.61\pm 0.36$ $2.82\pm 0.39$ $2.67\pm 0.33$ $23/30$ $22/30$ $26/30$ $78.89\%$
Ours $2.25\pm 0.23$ $2.39\pm 0.28$ $2.13\pm 0.24$ $28/30$ $/30$ 30 $29/30$ 96.67%
@@ V-D Manipulation Results
To evaluate the performance of the proposed SOI-based dynamics model for deformable fabric bag manipulation, we conducted SOI preserving and servoing experiments on a fabric bag with different oval shapes and the transitions between them. In SOI preserving experiments, we mainly compare the performance of our proposed approach with two commonly-used and well-established dynamics modeling techniques: the Mass-Spring Model (MSM) [35] and the Finite Element Model (FEM) [36]. The comparative analysis encompasses both quantitative and qualitative outcomes, as encapsulated in Figure 7(a) and Table III. Our findings reveal a discernible trend across all examined methods; there is an incremental rise in error corresponding to the transformation of the bag’s shape from a Long Oval (LO) to a Short Oval (SO), suggesting that as the bag opening rim becomes shorter and wider, the complexity of preserving its structure increases. Notably, our GNN-based dynamics model consistently outperforms the MSM and FEM across the entire spectrum of object shapes, achieving the lowest error rates in tasks dedicated to shape preservation. This outcome corroborates the superior expressive capability of our GNN-based model in capturing the intricate deformations of the fabric bags, and maintaining the SOI shape unchanged during the fabric bag moving, surpassing the traditional MSM and FEM approaches, particularly when dealing with objects that exhibit more complex dynamic behaviors. The advantage of our model is most pronounced when interacting with simpler geometric shapes.
In our SOI servoing experiments, we conduct a comprehensive analysis of the manipulation results associated with transitions between distinct shape categories of the structure of interest (SOI). Additionally, we juxtapose our dynamics model with two sophisticated shape servoing techniques tailored for deformable objects–visual servoing (VS) as described by Lagneau et al. [33], and the latent shape control (LSC) model introduced by Qi et al. [34]. The analysis integrates both quantitative and qualitative assessments, as depicted in Fig. 7(b) and detailed in Table IV. We consider three specific shape servoing tasks: LO (Long Oval) $\rightarrow$ SO (Short Oval), SO $\rightarrow$ RO (Round Oval), and LO $\rightarrow$ RO. The evaluation metrics include the Mean Hybrid Distance Error (recorded solely for successful trials) and the Success Rate of SOI servoing. Furthermore, we report the success rate for achieving complete servoing with a threshold error of less than $5$ cm, based on 30 trials for each method and shape transition scenario. The experimental outcomes reveal that our proposed method consistently secured the least shape error, ranging from $2.13$ to $2.39$ cm across all tested cases, and achieved an outstanding overall success rate of $96.67\%$. When contrasted with traditional model-based techniques such as the Mass-Spring Model (MSM) and the Finite Element Model (FEM), our data-driven SOI dynamics model exhibited a markedly enhanced capability in manipulating the SOI for the fabric bag. It was observed that transformations involving shorter and wider shapes incurred higher errors due to their inherently complex dynamics. Nevertheless, our proposed method demonstrated a robust performance across all tasks. In conclusion, the bimanual manipulation experiments underscore the effectiveness of the proposed SOI-based dynamics modeling approach, which ensures precise and dependable control over the shapes of deformable objects through advanced predictive modeling and optimization techniques.
@@ VI Conclusion
In this work, we introduced a novel bimanual manipulation framework for deformable bags, centered around a Structure of Interest (SOI)-based latent dynamics model. Our approach effectively integrates multi-view perception, graph neural networks, and model predictive control, facilitating precise and efficient robotic manipulation of flexible materials. Through simulations and real-world experiments, the framework demonstrated promising results in intelligent physical interaction with deformable objects. One limitation of the current system is its reliance on differently colored SOIs to distinguish target manipulation areas, which may not be feasible in all operational settings. Future efforts will aim to enhance the system’s ability to identify and manipulate SOIs without such visual aids, broadening the applicability of our method to a wider array of real-world scenarios.

### v2

@@ V-A Experiment Setup
Figure 7: Experimental set-up to validate bimanual deformable bag manipulation tasks. (a) Overview of the set-up. (b)-(d) The different SOI shape categories in bag experiments: Long Oval, Round Oval, and Short Oval. (e)-(g) The considered bag SOI and two additional types of T-shirt SOIs: opening rim, bottom hem and collar.
Figure 8: Results of our proposed global particle sampling approach. (Top) Reconstructions of different SOIs for different deformable objects. (Bottom) Corresponding SOI particle sets. (a)-(c): bag opening rim, (d)-(e): T-shirt bottom hem, and (f)-(g): T-shirt collar.
As illustrated in Fig. 7(b)-(d), our experiment considers three SOI shape categories for bag manipulation tasks: Long Oval (LO), Round Oval (RO), and Short Oval (SO). Additionally, we have incorporated experiments involving two additional types of T-shirt SOIs, each made from distinct fabric materials to further validate the effectiveness of our proposed approach on a variety of deformable objects. One SOI is the collar crafted from a soft modal fabric (Fig. 7(g)), while the other hem SOI comprises a slightly stiffer polyester-cotton blend (Fig. 7(f)). These variations in material properties provide a more comprehensive assessment of our method’s adaptability and robustness across different fabric types. For each deformable object, we collect a training dataset of 72,000 frames around 120 minutes, featuring 600 episodes and each spanning 120 frames. Within every episode, we execute 20 actions on the deformable objects. The data collection policy randomly selects from the action space parameters after doing collision checking and bag constraint checking, which include translations along the $x$, $y$, $z$ axes, and the rotation around $r_{z}$. Throughout each episode, we save the generated particle data with proposed sampling approach, as well as the robot end-effector’s poses.
The experiment comprises two types of manipulation tasks. In the SOI shape preservation task, the robotic manipulators aim to keep the SOI shape consistent while moving the fabric bag from its initial shape to a predefined target shape. Both the initial and target configurations share the same shape but differ in their spatial transformations. Conversely, the SOI shape servoing task involves changing the SOI shape from one configuration to another distinct shape configuration by manipulating the SOIs. We conduct a quantitative evaluation for each component of our proposed approach, including SOI particle sampling, the SOI-based neural dynamics model, and performance results of designed manipulation tasks. Our primary metrics for evaluating the manipulation performance are Chamfer distance (CD) and Earth mover’s distance (EMD), along with a hybrid metric that combines both. Furthermore, we employ Geodesic distance (GD) to assess the proximity of boundary points residing on one manifold with boundary.
TABLE II: Averaged SOI particle sampling results across 90 frames.
@@
V-B SOI Particle Sampling
We commence by benchmarking the proposed global particle sampling (GPS) technique against the local particle sampling (LPS) baseline. LPS initiates by processing the SOI-related partial point cloud of the deformable objects through its color filtering. Subsequently, it encapsulates the incomplete SOI point clouds with different convex hulls. Following this, it performs the point sampling procedure, eventually combining the sampled points into a full set of SOI particles. As shown in Table II, we compute the mean distance between the sampled particles and the ground-truth particles, the latter captured by a professional 3D scanner. Our analysis indicates that the GPS method incurs lower losses in terms of all distance metrics, thus outperforming the LPS approach. These outcomes align with the premise that incorporating additional topological information can markedly enhance the quality of sampling, particularly in scenarios where occlusions are present, and Fig. 8 presents example results for different SOIs applied to three selected deformable objects using our proposed GSP method.
Figure 9: (Left) SOI-based latent dynamics model training loss with different loss functions in testing data. (Right) The dynamics model’s evaluation over the time horizon in testing data.
Figure 10: (Left) Mean hybrid distance error and (Right) computation time over different control horizons and gradient steps in the MPC process. The white boundary represents the MPC computation time with respect to the horizon and gradient step being less than 20 ms, which is acceptable for the fabric manipulation task. Inside the boundary, the computation time for the specific setting of an MPC horizon of 6 and a gradient step of 11 is the most efficient.
V-C GNN-based SOI Dynamics Model
The training of our GNN model, detailed in Section IV-B, commences with the construction of a graph, where edges are formed between vertices that are within a distance threshold of $d=0.04$. By adopting a 70/15/15 split, we ensure that the GNN model has enough data to learn effectively while also enabling thorough validation and testing. Every vertex and edge are encoded using 3-layer Multilayer Perceptrons (MLPs), featuring hidden and output layers, each with 300 neurons. The propagation module is constituted by a fully connected neural layer with a layer size of 300. For motion prediction, we employ an additional 3-layer MLP with the hidden layer configured to 300 neurons. ReLU activation functions are utilized throughout the neural networks to introduce non-linearity. The model undergoes training for 120 epochs, employing the Adam optimizer. We have selected a batch size of 32 to balance the trade-off between generalization and computational efficiency. The learning rate is set at $5\mathrm{e}$-4, which is a conventional choice for steady convergence. These hyperparameters were chosen to foster a robust learning process while maintaining the capacity to capture complex patterns within the data. In practice, the final weights for the GNN-based SOI neural dynamics model are chosen based on the lowest validation loss recorded during training. This ensures that the model achieves the best possible performance on new, unseen data, which is critical for the practical application of the model in the integrated MPC for the bag manipulation task.
Subsequently, we evaluate the performance of GNN-based SOI dynamics models using different loss functions in testing data. Table III and Fig. 9(left) indicate that leveraging a hybrid loss function that integrates CD and EMD leads to superior performance across all metrics, as opposed to optimizing individually for either CD or EMD. Notably, the Geodesic distance is significantly improved, compared to using them individually. The combination of $0.15$ CD + $0.85$ EMD achieves the best performance across all metrics, giving the lowest CD, competitive EMD, and lowest GD, so we select this hybrid distance for the following experiments. In Fig. 9(right), the dynamics model is evaluated on testing data using the CD and the EMD loss over the time horizon. Our analysis reveals that while an unavoidable increase in loss metrics occurs as the prediction horizon of the model is extended, these metrics values remain within an acceptable range. This outcome validates the potential utility of this model in following deformable object manipulation tasks, demonstrating their capacity to maintain reasonable accuracy over a longer prediction horizon.
To evaluate the performance of the proposed SOI-based neural dynamics model for real deformable object manipulation tasks, we integrate it with MPC to conduct SOI-preserving experiments on a fabric bag with various oval shapes. To balance real-time performance and control accuracy for our manipulation task, we set the MPC horizon to 6 and the gradient step to 11, based on extensive empirical testing (see Fig. 10). This combination provides a reasonable compromise between computational demand and the quality of optimization for the MPC. We compare the performance of our proposed approach with two commonly-used and well-established dynamics modeling techniques: the Mass-Spring Model (MSM) [7] and the Finite Element Model (FEM) [8]. These two molding techniques are also integrated into the MPC framework for conducting the experiments. The comparative analysis over 60 trails includes both quantitative and qualitative outcomes, as shown in Table IV, Fig. 11(a) and Fig. 12(a). Our findings reveal a clear trend across all examined methods; there is an incremental rise in error corresponding to the transformation of the bag’s shape from a Long Oval (LO) to a Short Oval (SO), suggesting that as the bag opening rim becomes shorter and wider, the complexity of preserving its structure increases. Notably, our GNN-based dynamics model consistently outperforms the MSM and FEM across all object shapes, achieving the lowest hybrid loss in tasks dedicated to shape preservation. This outcome indicates the superior expressive capability of our GNN-based model in capturing the intricate deformations of the fabric bags, and maintaining the SOI shape unchanged during the fabric bag moving, surpassing the traditional MSM and FEM approaches, particularly when dealing with objects that exhibit more complex dynamic behaviors. The advantage of our model is most pronounced when interacting with simpler geometric shapes.
@@ V-C GNN-based SOI Dynamics Model
TABLE III: The performance of the SOI particle dynamics model with different loss functions in testing data
@@
TABLE IV: Mean Hybrid Distance Error for SOI Shape Preserving by Different Dynamics Modeling Methods over 60 Trials
@@
TABLE V: Mean Hybrid Distance Error and Success Rate for SOI Shape Preserving on Different Deformable Objects by Different Methods
Method Bag Opening Rim (180 trls) Soft T-shirt Collar (60 trls) Stiff T-shirt Bottom Hem (60 trls) Total
Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$
VS [30] $3.20\pm 0.20$ $71.11\%$ $3.41\pm 0.21$ $76.67\%$ $3.28\pm 0.29$ $68.33\%$ $71.67\%$
MSM [7] $2.80\pm 0.37$ $73.89\%$ $2.48\pm 0.40$ $71.67\%$ $3.06\pm 0.48$ $65.00\%$ $71.67\%$
FEM [8] $2.17\pm 0.34$ $83.33\%$ $2.23\pm 0.33$ $80.00\%$ $2.64\pm 0.38$ $86.67\%$ $83.33\%$
G-DOOM [24] $3.07\pm 0.32$ $79.44\%$ $2.64\pm 0.34$ $75.00\%$ $3.31\pm 0.43$ $71.67\%$ $77.00\%$
Ours $1.86\pm 0.23$ $\mathbf{98.33\%}$ $1.65\pm 0.27$ $\mathbf{95.00\%}$ $2.43\pm 0.26$ $\mathbf{86.67\%}$ $\mathbf{95.33\%}$
TABLE VI: Mean Hybrid Distance Error and Success Rate for SOI Shape Servoing on Different Deformable Objects by Different Methods
Method Bag Opening Rim (180 trls) Soft T-shirt Collar (60 trls) Stiff T-shirt Bottom Hem (60 trls) Total
Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$ Mean HD Error $\downarrow$ Success Rate $\uparrow$
VS [30] $3.76\pm 0.21$ $68.89\%$ $3.75\pm 0.19$ $71.67\%$ $3.84\pm 0.27$ $70.00\%$ $69.67\%$
MSM [7] $3.29\pm 0.45$ $70.00\%$ $2.94\pm 0.42$ $66.67\%$ $3.57\pm 0.47$ $63.33\%$ $68.00\%$
FEM [8] $2.70\pm 0.35$ $78.89\%$ $2.62\pm 0.37$ $76.67\%$ $2.88\pm 0.39$ $80.00\%$ $78.67\%$
G-DOOM [24] $3.48\pm 0.33$ $75.56\%$ $3.11\pm 0.37$ $68.33\%$ $3.87\pm 0.41$ $66.67\%$ $72.33\%$
Ours $2.27\pm 0.24$ $\mathbf{96.67\%}$ $1.87\pm 0.26$ $\mathbf{91.67\%}$ $2.68\pm 0.29$ $\mathbf{90.00\%}$ $\mathbf{94.33\%}$
Figure 11: Qualitative results in (a) SOI shape preserving and (b) SOI shape servoing tasks with our proposed approach.
Figure 12: Achieved SOI particle configurations of (a) SOI shape preserving and (b) SOI shape servoing tasks by different approaches.
@@ V-D Manipulation Results
[…]
@@ VI Conclusion
In this work, we introduced a novel bimanual manipulation framework for deformable bags, centered around a SOI-based latent dynamics model. Our approach effectively integrates multi-view perception, graph neural networks, and model predictive control, facilitating precise and efficient robotic manipulation of flexible materials. Through real-world experiments, the framework demonstrated promising results in intelligent physical interaction with deformable objects. One limitation of the current system is its reliance on differently colored SOIs to distinguish target manipulation areas, which may not be feasible in all operational settings. Future efforts will focus on employing a learning-based SOI detection module and multimodal sensing to enhance the system’s ability to identify and manipulate SOIs without such visual aids, thereby broadening the applicability of our method to a wider array of real-world scenarios.